OpenStudy (anonymous):

Suppose you are playing a game with two number cubes. Let A represent rolling 2, 3, or 4, and B represent rolling 1, 5, or 6. The probability of A is 1/2 and the probability of B is 1/2 . a. Simplify (1/2 A + 1/2 B)^2 b. What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6?

OpenStudy (anonymous):

i don't mean to be a critic, but this line \((\frac{1}{2}A+\frac{1}{2}B)^2\) makes no sense in this context, because A and B are sets, not numbers

OpenStudy (anonymous):

answer to part two is \(\frac{1}{2}\times \frac{1}{2}\)

OpenStudy (anonymous):

that is exactly how it is written just how you posted it

OpenStudy (anonymous):

i am confused on A.)

OpenStudy (anonymous):

yeah me too, it makes no sense whatsoever

OpenStudy (anonymous):

A is a set B is a set you cannot multiply a set by \(\frac{1}{2}\) and add it to another set. it make no sense

OpenStudy (anonymous):

you add and multiply numbers, not sets you can take the union of two sets if you like, or their intersection, but you cannot add or multiply

OpenStudy (anonymous):

where did the question come from ?

OpenStudy (anonymous):

a test

OpenStudy (anonymous):

and that is exactly how they put it on there the same way i posted it and you rewrote it the first time how the problem actually looks on the test

OpenStudy (anonymous):

well it make no sense if A and B were numbers or variables that represent numbers you could do it. if they are sets (it says " Let A represent rolling 2, 3, or 4") then you cannot do it. sorry

OpenStudy (anonymous):

you might want to ask your teacher what he or she had in mind, and ask (gently) how one is expected to multiply sets together. i take it this is not in a text book (i hope it is not)

OpenStudy (anonymous):

Thank you for the help